if x = 3 + 2 root 2 then value of root x + 1/ root x =

Grade:9

2 Answers

Akshay
8 years ago
3+rt2 = (1+rt2)2, hope that will help
ATB
Manoj
5 years ago
1/x=1/3+^2*3-2^2/3-2^2=3-2^2/(3)^2-(2^2)^2=3-2^2/9-8=3-2^2.                                                             
(^x-1/^x)^2=x-1/x+2=(3+2^2)+(3-2^2)+2=8.                                                                                                  
So,(^x-1/^x)^2=^8-2^2=2^2